A lightning bolt strikes 30 m from a circular metal loop of area 0.5 m^2. As the lightning bolt current rises from zero to 6 x 10^6 amps in 10 microseconds, what is the magnitude of the average induced emf in the loop?


Please explain the process using non-calculus based steps.
Answer #1

B = mu*I/2*pi*R

V = d(Fi)/dt

=Fi = B*A

=2*10^-7*6*10^6/30*0.5

= 0.02

V = d(fi)/dt

= 0.02/0.01

= 2 V